solve the following system of equations
x+5y= 8 (1)
x= 5-5y (2)

Answer by stanbon(75887)

solve the following system of equations
x+5y= 8
x= 5-5y
-----------
Rearrange:
x + 5y = 8
x + 5y = 5
-------
Assuming those x's and y's mean the same
in both equations you subtract and get:
---
0 = 3
--
The assumption was wrong. There is no (x,y)
pair that satisfies both equations.
---
Conclusion: No solution
==============================
Cheers,
Stan H.